OpenStudy (anonymous):

Molar Mass (g/mol) HNO3 63.013 N atom 14.007 H atom 1.0080 O atom 15.999 Avogadro's No. (mol-1) 6.022×1023

OpenStudy (anonymous):

If a sample of HNO3 is found to contain 6.74 · 1024 N atoms how many moles of H atoms are present?

OpenStudy (anonymous):

n=N/L = 6,74 *10^24 / 6,022*10^23 mol-1= 11,19 mol of "N" and ratio of N and H in HNO3 is same so n(H) = 11,19 mol